# Calibrating Translation Decoding with Quality Estimation on LLMs

### v1

@@ Abstract
Neural machine translation (NMT) systems typically employ maximum a posteriori (MAP) decoding to select the highest-scoring translation from the distribution mass. However, recent evidence highlights the inadequacy of MAP decoding, often resulting in low-quality or even pathological hypotheses — the decoding objective is not aligned with real-world translation quality. This paper proposes calibrating hypothesis likelihood with translation quality from a distributional view by directly optimizing their Pearson correlation — thereby enhancing the effectiveness of translation decoding. With our method, translation on large language models (LLMs) improves substantially after limited training (2K instances per direction). This improvement is orthogonal to those achieved through supervised fine-tuning, leading to substantial gains across a broad range of metrics and human evaluations — even when applied to top-performing translation-specialized LLMs fine-tuned on high-quality translation data, such as Tower, or when compared to recent preference optimization methods, like CPO. Moreover, the calibrated translation likelihood can directly serve as a strong proxy for translation quality, closely approximating or even surpassing some state-of-the-art translation quality estimation models, like CometKiwi. Lastly, our in-depth analysis demonstrates that calibration enhances the effectiveness of MAP decoding, thereby enabling greater efficiency in real-world deployment. The resulting state-of-the-art translation model, which covers 10 languages, along with the accompanying code and human evaluation data, has been released to the community: https://github.com/moore3930/calibrating-llm-mt.
@@ 1 Introduction
Prior studies have tried to mitigate this miscalibration issue by introducing an additional optimization step during inference time, as shown in Figure 1 upside, known as quality-aware decoding (QAD) (Fernandes et al., 2022). These approaches typically involve generating multiple candidate translations through sampling, followed by reranking or voting using reference-free and/or reference-based machine translation metrics, such as Best-of-N (BoN) sampling (Rei et al., 2024a; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a). While effective, this line of methods inevitably incurs significantly higher latency. For example, the current top-performing system in the WMT24 competition (Kocmi et al., 2024a), Tower (Rei et al., 2024a), requires sampling 100 times for each prompt from a 70B LLM-based translation model; this is impractical, particularly for online environments.
This paper proposes a simple yet highly effective method to directly optimize the calibration during training, where we encourage the likelihoods of translation hypotheses to align more closely with their quality distribution by directly optimizing Pearson correlations between them. More specifically, as shown in Figure 1 lower side, given a translation prompt $x$, we sample multiple hypotheses from an NMT model during training time and compute both their likelihoods and quality scores, as measured by any external metric, such as COMET. We then define the loss as the negative Pearson correlation between the two sets of data points and minimize it directly using a standard gradient-based optimizer. While simple, our approach yields substantial performance improvements with limited training — using only 2K instances per language — even when applied to state-of-the-art LLM-based MT systems such as Tower. Moreover, due to clear calibration improvements, the resulting models’ likelihood can directly serve as a strong proxy for translation quality, even surpassing some state-of-the-art translation quality estimation (QE) models, like Comet-Kiwi (Rei et al., 2022a).
Figure 1: The illustration of quality-aware decoding (upper side) and our calibration method (lower side). The former explores the decoding space by adding an extra step during test-time decoding, which involves multiple rounds of sampling followed by reranking or voting to select the highest-scoring output. Our method focuses on training time optimization, intending to calibrate the distribution of hypothesis likelihoods to the corresponding qualities.
@@ 2 Background
At a practical level, this study aims to: (1) optimize translation quality, and (2) calibrate the model’s uncertainty quantification, i.e., likelihood, with human judgments. To this end, we briefly introduce key foundational concepts, including metric meta-evaluation and uncertainty in translation.
@@
Translation uncertainty. Uncertainty estimation (Ott et al., 2018; Huang et al., 2024) is an important diagnostic tool often used to assess confidence in the model’s output. Previous work (Galley et al., 2015; Fomicheva et al., 2020) has shown that, under a Bayesian framework, appropriate quantification, such as entropies, posterior distributions, or attention weights, can yield uncertainty estimates that correlate well with human judgments of translation quality, even without human-labeled training data. In this paper, we focus on a specific form of quantification — likelihood — due to its natural connection to translation training and decoding.
Previous approaches.
A few different themes investigate optimizing translation performance using signals from downstream quality evaluation metrics. We briefly summarize as follows.
  1. 1.
    Test-time quality optimization. To alleviate the gap between decoding objective and translation quality, Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a; Yang et al., 2024) suggests basing decoding decisions on statistics gathered from the distribution holistically. As reference-free quality estimation (QE) progresses, Best-of-N sampling (Rei et al., 2024b; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) with a QE model becomes a straightforward strategy. However, both of these methods will result in a significant amount of additional decoding time. Moreover, the risk of metric bias increases, as translation improvements may stem from metric hacking rather than genuine quality enhancements (Skalse et al., 2022; Kovacs et al., 2024).
  2. 2.
    Preference optimization. Recently, direct preference optimization (DPO) and its variants (Rafailov et al., 2023; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024) have emerged as promising approaches for post-training LLMs, providing efficient alternatives to reinforcement learning from human feedback (RLHF). This line of methods often employs pairwise preferences under the Bradley-Terry framework (Bradley and Terry, 1952) to provide rewards, or a more general Plackett-Luce ranking framework (Plackett, 1975) when multiple ranked hypotheses are available — a few studies (Xu et al., 2024; Zhu et al., 2024) in translation also follow them. Additionally, some works (Guo et al., 2024; Lambert et al., 2024) employ these methods in an on-policy training paradigm for online LLMs alignments. Notably, regardless of employing pairwise/listwise data or on-/off-policy training, the objective is generally to maximize expected rewards. One key distinction of our method, as will be shown in §3, is that our training objective is to maximize correlation instead of expected rewards. Therefore, we follow prior research (Ott et al., 2018; Zhao et al., 2022) and use the term "calibration" in this paper to highlight the difference. More detailed discrepancies can be found in the next section.
@@ 4.1 Base Models
We base our experiments on two series of strong translation-specified LLMs, i.e., ALMA-Base (Xu et al., 2023) and Tower-Base (Alves et al., 2024), in both 7B and 13B variants. Both models are obtained by continued pretraining of LLAMA2-7B and LLAMA2-13B on extensive multilingual data. A key distinction is that ALMA-Base relies exclusively on monolingual data (Xu et al., 2023) for continued pretraining, whereas Tower-Base incorporates both parallel and monolingual data to improve cross-lingual capabilities. Both models achieve remarkable translation performance. For instance, the Tower series systems (Rei et al., 2024b) won all of the translation tasks in the WMT24 general translation tracks (Kocmi et al., 2024a), showing state-of-the-art performance.
@@ 4.2 Data and Evaluation
Translation evaluation dataset. For fair comparison with the ALMA and Tower model series, we evaluate translation performance on the WMT22 and WMT24 datasets (Zerva et al., 2022; Kocmi et al., 2024a), respectively. Detailed dataset descriptions can be found in Appendix A. Following current best practice (Freitag et al., 2022b, 2023), all of the results in this paper are measured by current widely used neural metrics, involving reference-based metric COMET, and the reference-free metrics XCOMET, CometKiwi-XL, and CometKiwi-XXL [^1] The corresponding metric model versions are Unbabel/wmt22-comet-da, Unbabel/XCOMET-XXL, Unbabel/wmt23-cometkiwi-da-xl, and Unbabel/wmt23-cometkiwi-da-xxl, respectively..
@@ 5.1 Calibration Leads to Clear Quality Improvements
We also conducted supervised fine-tuning on the TowerBase series using 2K Best-of-N samples per direction, selected from our calibration dataset (see §4.2) based on the highest CometKiwi-XXL scores. It is easy to see that the resulting average performances are comparable to those of official Instruction models, highlighting the quality of our dataset. Note that performance degrades when fine-tuning on the full calibration set, as some of the sampled bitext examples are of low quality.
@@
We also compare with CPO (Xu et al., 2024), a widely used preference optimization method in translation, following their original setting and selecting the highest- and lowest-scoring sampled candidates as accepted and rejected samples, respectively — consistent and substantial improvements are observed. Additionally, Appendix B.2 reports results based on the ALMA base model, showing gains of +2.0 XCOMET and +1.4 KIWI-XXL for out-of-English translation over CPO.
@@ 5.2 Calibration Leads to Direct Quality Estimation
We use the WMT22 metric meta-evaluation dataset (Zerva et al., 2022) and follow the official practice, using the multidimensional quality metrics (MQM) scores as golden quality annotations, which are provided by expert translators and are considered more reliable than other annotations. We evaluate all training directions on Tower that overlap with the WMT dataset, namely, ende and enru.
@@
  • The TowerInstruct-7B/-13B models exhibit relatively low performance in both the translation and quality estimation abilities. The resulting models manifest progressively improved translation performance as more best-of-N samples are included for supervised fine-tuning. Interestingly, the corresponding quality estimation abilities, measured by Spearman scores calculated on the WMT22 dataset, increase from around 51.5 to 54.0 points. We attribute this to fine-tuning on Best-of-N samples, which are predominantly high-quality, encouraging the model to assign higher likelihoods to better hypotheses.
  • Examining the effects of scaling, we observe that: (1) scaling up from 7B to 13B generally improves translation performance for both the original TowerInstruct models and the fine-tuned models using our best-of-N data; (2) however, its impact on calibration — that is, the models’ quality estimation ability — remains minimal.
  • Our calibration method manifests very strong improvements in both translation and quality estimation. For example, when applying our method to TowerBase-13B, the resulting model surpasses some state-of-the-art systems in both translation performance and quality estimation ability, i.e., Tower-70B-v2+MBR/TRR and CometKiwi, simultaneously.
@@ 5.3 Calibration Enhances Effectiveness for Maximum a Posteriori Decoding
Figure 3: The average performance (ende,es,ru,zh) measured by COMET score for both TowerInstruct-Mistral-7B and our calibrated models, when (1) varying sampling size for Best-of-N sampling and (2) applying beam search with a beam size of 5.
@@ 7 Conclusion
Very clear and consistent performance improvements.
Even when applied to state-of-the-art systems such as Tower, our calibration method yields substantial improvements across a wide range of automatic metrics and human evaluations. For example, a calibrated Tower-7B model with beam search achieves state-of-the-art translation performance comparable to — if not exceeding — that of the much larger Tower-70B-v2 model (which employs test-time optimizations such as MBR decoding), while the calibrated Tower-7B model requires substantially less computation.
A unified view of quality optimization and estimation in translation.
Our approach adopts a shared objective for both quality optimization and quality estimation in translation, offering a unified perspective on these two tasks. Our empirical analysis also supports this perspective, indicating a clear correlation between calibration level and translation performance. Experimentally, the calibrated method achieves both state-of-the-art translation performance and accurate quality estimation within a single model, with the latter even surpassing widely used QE models such as CometKiwi — all without relying on human-annotated data.
Enhanced realization for maximum a posterior decoding.
By calibrating the training objective with real-world quality, the effectiveness of widely used efficient MAP decoding rules, such as beam search, is improved. This paper shows that the effectiveness of beam search originally fell far short of Best-of-N sampling in translation quality for the baseline system. However, our method not only substantially improves performance, but also enables beam search to achieve comparable performance to Best-of-N sampling with 100 hypotheses using only a beam size of 5 — dramatically reducing inference cost without sacrificing quality, showing efficiency potential for real-world deployment.

### v2

@@ Abstract
Neural machine translation (NMT) systems typically employ maximum a posteriori (MAP) decoding to select the highest-scoring translation from the distribution. However, recent evidence highlights the inadequacy of MAP decoding, often resulting in low-quality or even pathological hypotheses as the decoding objective is only weakly aligned with real-world translation quality. This paper proposes to directly calibrate hypothesis likelihood with translation quality from a distributional view by directly optimizing their Pearson correlation, thereby enhancing the effectiveness of translation decoding. With our method, translation with large language models (LLMs) improves substantially after limited training (2K instances per direction). This improvement is orthogonal to those achieved through supervised fine-tuning, leading to substantial gains across a broad range of metrics and human evaluations. This holds even when applied to top-performing translation-specialized LLMs fine-tuned on high-quality translation data, such as Tower, or when compared to recent preference optimization methods, like CPO. Moreover, the calibrated translation likelihood can directly serve as a strong proxy for translation quality, closely approximating or even surpassing some state-of-the-art translation quality estimation models, like CometKiwi. Lastly, our in-depth analysis demonstrates that calibration enhances the effectiveness of MAP decoding, thereby enabling greater efficiency in real-world deployment. The resulting state-of-the-art translation model, which covers 10 languages, along with the accompanying code and human evaluation data, has been released: https://github.com/moore3930/calibrating-llm-mt.
@@ 1 Introduction
Prior studies have tried to mitigate this miscalibration issue by introducing an additional optimization step during inference time, as shown in Figure 1 (top), known as quality-aware decoding (QAD) (Fernandes et al., 2022). These approaches typically involve generating multiple candidate translations through sampling, followed by reranking or voting using reference-free and/or reference-based machine translation metrics, such as Best-of-N (BoN) sampling (Rei et al., 2024a; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a). While effective, this line of methods inevitably incurs significantly higher latency. For example, the current top-performing system in the WMT24 competition (Kocmi et al., 2024a), Tower (Rei et al., 2024a), requires sampling 100 times for each prompt from a 70B LLM-based translation model; this is impractical, particularly for online environments.
This paper proposes a simple yet highly effective method to directly optimize the calibration during training, where we encourage the likelihood of translation hypotheses to align more closely with their quality distribution by directly optimizing Pearson correlations between them. More specifically, as shown in Figure 1 (bottom), given a translation prompt $x$, we sample multiple hypotheses from an NMT model during training time and compute both their likelihoods and quality scores, as measured by any external metric, such as COMET. We then define the loss as the negative Pearson correlation between the two sets of data points and minimize it directly using a standard gradient-based optimizer. While simple, our approach yields substantial performance improvements with limited training — using only 2K instances per language — even when applied to state-of-the-art LLM-based MT systems such as Tower. Moreover, due to clear calibration improvements, the resulting models’ likelihood can directly serve as a strong proxy for translation quality, even surpassing some state-of-the-art translation quality estimation (QE) models, like Comet-Kiwi (Rei et al., 2022a).
Figure 1: The illustration of quality-aware decoding (top) and our calibration method (bottom). The former explores the decoding space by adding an extra step during test-time decoding, which involves multiple rounds of sampling followed by reranking or voting to select the highest-scoring output, incurring high costs. Our method focuses on training-time optimization, aiming to calibrate the likelihoods of hypotheses to their corresponding quality scores, enabling effective MAP decoding.
@@ 2 Background
At a practical level, this study aims to optimize: (1) translation quality and (2) quality estimation, simultaneously in a single model. To this end, we briefly introduce related key foundational concepts.
@@
Translation quality optimization.
A few different themes investigate optimizing translation performance using signals from external metric models. We briefly summarize as follows:
(1) Test-time quality optimization. To alleviate the gap between decoding objective and translation quality, Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a; Yang et al., 2024) suggests basing decoding decisions on statistics gathered from the distribution holistically. As reference-free quality estimation (QE) progresses, Best-of-N sampling (Rei et al., 2024b; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) with a QE model becomes a straightforward strategy. However, both of these methods will result in a significant amount of additional decoding time. Moreover, the risk of metric bias increases, as translation improvements may stem from metric hacking rather than genuine quality enhancements (Skalse et al., 2022; Kovacs et al., 2024).
(2) Preference optimization. Recently, direct preference optimization (DPO) and its variants (Rafailov et al., 2023; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024) have emerged as promising approaches for post-training LLMs, providing efficient alternatives to reinforcement learning from human feedback (RLHF). This line of methods often employs pairwise preferences under the Bradley-Terry framework (Bradley and Terry, 1952) to provide rewards, or a more general Plackett-Luce ranking framework (Plackett, 1975) when multiple ranked hypotheses are available — a few studies (Xu et al., 2024; Zhu et al., 2024) in translation also follow them. Additionally, some works (Guo et al., 2024; Lambert et al., 2024) employ these methods in an on-policy training paradigm for online LLMs alignments. Notably, regardless of employing pairwise/listwise data or on-/off-policy training, the objective is generally to maximize expected rewards. One key distinction of our method, as will be shown in §3, is that our training objective is to maximize correlation instead of expected rewards. Therefore, we follow prior research (Ott et al., 2018; Zhao et al., 2022) and use the term "calibration" in this paper to highlight the difference. More detailed discrepancies can be found in the next section.
@@ 4.1 Base Models
We base our experiments on two series of strong translation-specified LLMs, i.e., ALMA-Base (Xu et al., 2023) and Tower-Base (Alves et al., 2024), in both 7B and 13B variants. Both models achieve remarkable translation performance after post-training. For instance, the Tower series systems (Rei et al., 2024b) achieved state-of-the-art performance by winning all translation tasks in the WMT24 general translation tracks (Kocmi et al., 2024a), outperforming both Google Translate and GPT-4.
@@ 4.2 Data and Evaluation
Translation evaluation dataset. For fair comparison with the ALMA and Tower model series, we evaluate translation performance on the WMT22 and WMT24 datasets (Zerva et al., 2022; Kocmi et al., 2024a), respectively. Detailed dataset descriptions can be found in Appendix A. Following current best practice (Freitag et al., 2022b, 2023), all of the results in this paper are measured by widely used neural metrics, involving reference-based metric COMET, and the reference-free metrics XCOMET, CometKiwi-XL, and CometKiwi-XXL [^1] The corresponding metric model versions are Unbabel/wmt22-comet-da, Unbabel/XCOMET-XXL, Unbabel/wmt23-cometkiwi-da-xl, and Unbabel/wmt23-cometkiwi-da-xxl, respectively..
@@ 5.1 Calibration Leads to Clear Quality Improvements
We also conducted supervised fine-tuning on the TowerBase series using 2K Best-of-N samples per direction, selected from our calibration dataset (see §4.2) based on the highest CometKiwi-XXL scores. The resulting average performances are comparable to those of official Instruction models. Note that the performance degradation when fine-tuning on the full calibration set is expected, as some of the sampled bitext examples are of low quality.
@@
We also compare with CPO (Xu et al., 2024), a widely used preference optimization method in translation, following its original setting by selecting the highest- and lowest-scoring sampled candidates as accepted and rejected samples, respectively — consistent and substantial improvements over CPO are observed. Additionally, Appendix B.2 reports results based on the ALMA base model, showing gains of +2.0 XCOMET and +1.4 KIWI-XXL for out-of-English translation over CPO.
@@ 5.2 Calibration Leads to Direct Quality Estimation
We use the WMT22 metric meta-evaluation dataset (Zerva et al., 2022) and follow the official practice, see § 4.2, to assess quality estimation ability using Spearman’s and Kendall’s correlation. We evaluate all training directions on Tower that overlap with the WMT dataset, namely, ende and enru.
@@
(1) As more best-of-N samples are included in supervised fine-tuning, translation performance improves progressively. Interestingly, the corresponding quality estimation abilities, measured by Spearman scores, increase from around 51.5 to 54.0 points. We attribute this to the fact that fine-tuning on Best-of-N samples encourages the model to assign higher likelihoods to better hypotheses.
(2) Examining the effects of scaling, we observe that: (1) scaling up from 7B to 13B generally improves translation performance for both the original TowerInstruct models and the fine-tuned models; (2) however, its impact on calibration — quality estimation ability — remains minimal.
(3) Our calibration method manifests very strong improvements in both translation and quality estimation. For example, when applying our method to TowerBase-13B, the resulting model surpasses some state-of-the-art systems in both translation performance and quality estimation ability, i.e., Tower-70B-v2+MBR/TRR and CometKiwi, simultaneously.
@@ 5.3 Calibration Enhances Effectiveness for Maximum a Posteriori Decoding
Figure 3: The average performance (ende,es,ru,zh) measured by COMET score for both TowerInstruct-Mistral-7B and our calibrated models, when varying sampling size for Best-of-N (BoN) sampling and applying beam search with a beam size of 5.
@@ 7 Conclusion
Very clear and consistent performance improvements. Even when applied to state-of-the-art systems such as Tower, our calibration method yields substantial improvements across a wide range of automatic metrics and human evaluations. For example, a calibrated Tower-7B model with beam search achieves state-of-the-art translation performance comparable to — if not exceeding — that of the much larger Tower-70B-v2 model (which employs test-time optimizations such as MBR decoding), while the calibrated Tower-7B model requires substantially less computation.
A unified view of quality optimization and estimation in translation. Our approach adopts a shared objective for both quality optimization and quality estimation in translation, offering a unified perspective on these two tasks. Our empirical analysis also supports this perspective, indicating a clear correlation between calibration level and translation performance. Experimentally, the calibrated method achieves both state-of-the-art translation performance and accurate quality estimation within a single model, with the latter even surpassing widely used QE models such as CometKiwi — all without relying on human-annotated data.
Enhanced realization for maximum a posterior decoding. By calibrating the training objective with real-world quality, the effectiveness of widely used efficient MAP decoding rules, such as beam search, is improved. This paper shows that the effectiveness of beam search originally fell far short of Best-of-N sampling in translation quality for the baseline system. However, our method not only substantially improves performance, but also enables beam search to achieve comparable performance to Best-of-N sampling with 100 hypotheses using only a beam size of 5 — dramatically reducing inference cost without sacrificing quality, showing efficiency potential for real-world deployment.